I am trying to solve this problem by using the elimination method.
0.05x + 0.025y = 11
0.15x + 0.05y = 12
Thanks :o)

Answer by AAfter Search(61)

0.05x + 0.025y = 11 .... (i)
0.15x + 0.05y = 12 ..... (ii)
By (i) x 2, we get
0.10x + 0.050y = 22
Hence,
0.10x + 0.050y = 22
0.15x + 0.050y = 12
------------------
(subtracting) -0.05x = 10
=> x = 10/(-0.05) = -200
Putting the value of x in equation (i), we get
0.05(-200) + 0.025y = 11
=> -10 + 0.025y = 11
=> 0.025y = 11 + 10 = 21
=> y = 21/0.025 = 840